OpenStudy (anonymous):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (texaschic101):

Is this an arithmetic sequence or an arithmetic sequence ?

OpenStudy (texaschic101):

oops..or a geometric sequence

OpenStudy (anonymous):

These is an arithmetic sequence

OpenStudy (texaschic101):

In an arithmetic sequence, you find the number by adding. -10 + (-2) = -10 - 2 = -12 (1st term) -12 + (-2) = -12 - 2 = -14 (2nd term) -14 + (-2) = -14 - 2 = -16 (3rd term) you get the idea ?

OpenStudy (texaschic101):

oops..-12 is the second term....and so on...

OpenStudy (anonymous):

I think so I would just continue that pattern until the 10th term?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

So would the 10th term be -30?